A mono chromatic source of light operating at 200w emits 4×10^20 protons per second. Find the wave length of the light

Grade:12

1 Answers

Arun
5 years ago
 
The energy of each photon=
E = (200 J/sec)/(4 * 10^20 /sec)
= 5 * 10^-19 J
 
Now \lambda = hc/ E
\lambda = (6.6* 10^-34) *(3* 10^8)/(5 * 10^-19)
400 nano metre